OpenStudy (anonymous):

A 0.300 kg baseball is thrown in a straight line through the air. At a height of 4.5 m above the surface of the earth, it has speed of 15.0 m/s. what is the total mechanical energy of the baseball?

OpenStudy (anonymous):

\[\frac12mv^2 + mgh = .3\left(\frac12 15\cdot 15 + 4.5 \cdot 10\right)\]

OpenStudy (anonymous):

4050